# Geometry of quantum states and chaos-integrability transition

### v1

@@ VI.1 Fidelity susceptibility for tridiagonal Gaussian $\beta$-ensembles
One difference from the situation considered in the two previous sections is that here we are considering the simplest situation where the total Hamiltonian $H$ has only one parameter $r$, rather than both $r$ and $\phi$. Hence, in this case, the line element will consist of a single element, which is the fidelity susceptibility associated with the parameter $r$. The QMT components in such a situation are evaluated numerically in the next subsection.
For our purposes, it is sufficient to consider a system with a Hamiltonian of the form $H=H_{0}+r~{}\mathcal{H}_{\beta}~{}$ where $H_{0}$ is a diagonal ‘integrable’ Hamiltonian, i.e., we assume that its eigenvalues are drawn from independent Gaussian distributions with zero mean and variance $1$, and $\mathcal{H}_{\beta}~{}$ is a ‘tridiagonal’ matrix. Therefore, we take $\mathcal{H}_{\beta}$ to be the following $2\times 2$ matrix
${}~{}~{}\mathcal{H}_{\beta}=\left(\begin{array}[]{ccc}a_{0}&b_{1}\\ b_{1}&a_{1}\end{array}\right)~{}.$ (67)
@@
The Hamiltonian $H=H_{0}+r~{}\mathcal{H}_{\beta}~{}$ can be thought of as the generalisation of the usual RP model (which is usually defined with $\mathcal{H}$ taken from the classical Gaussian $\beta$-ensembles) for general values of $\beta$. (citations) The expression for $g_{rr}$, averaged over the ensemble is given by,
@@
Here $\mathcal{H}_{mn}=\braket{m}{\mathcal{H}_{\beta}}{n}$, are the matrix elements of $\mathcal{H}_{\beta}$ in the energy eigenbasis of the total Hamiltonian. Note that the matrix elements $\mathcal{H}_{12}$ and $\mathcal{H}_{21}$ are zero if for a realisation of the matrix $H_{\beta}$ the condition $h_{1}=h_{2}$ is satisfied.
@@ VI.2 QMT for the tridiagonal $\beta$-ensemble
So far in this section, we have considered the simplest situation where the total Hamiltonian contains one parameter and the matrices are two-dimensional. Here we now consider the more general situation where the total Hamiltonian is assumed to be of the form (54), where the $N\times N$ matrices $\mathcal{H}_{i}$ are now sampled from tridiagonal $\beta$-ensembles with generic Dyson indices. In Fig. 8 we have plotted the numerical values of the metric components for different fixed values of $r$ (and with a fixed $\phi$), when the matrix $H_{0}$ is drawn from a $\beta$-ensemble with a Dyson index $\gamma<1$, and $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are drawn from $\beta$-ensemble with the same Dyson index $\beta>1$. Note that, compared with the previous cases of GUE Hamiltonians, the mixed component $g_{r\phi}$ is non-zero, and the numerical values of this component are shown in the inset of Fig. 8.
@@
Figure 8: Log-Log (with base 10) plot of the components of the metric tensor for Gaussian $\beta$-ensemble with large $N$. The Hamiltonian is of the same form as (54), with the individual matrices being drawn from $\beta$-ensemble with different Dyson indices ($H_{0}$ is from ensemble with Dyson index $\gamma<1$, while $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are drawn from $\beta$-ensemble with the same Dyson index $\beta>1$). The red and the orange dots show the numerical values of the $g_{rr}$ and $g_{\phi\phi}$ components, respectively, for fixed values of $r$ and with $\beta=2$ and $\gamma=0.5$, while the brown and black dots represent these components when $\beta=3$ and $\gamma=0.2$. The plot in the inset shows the values of the $g_{r\phi}$ component for these two cases (magenta: $\beta=2$ and $\gamma=0.5$, and green: $\beta=3$ and $\gamma=0.2$). For all the plots, we have fixed $N=1000$, and the results show expressions averaged over 1000 different realisations of the ensemble.
@@ VII The fidelity-susceptibility for an invariant $\beta$-ensemble with an effective Dyson index
In the previous section we have obtained the FS associated with a $2\times 2$ random matrix ensemble with general Dyson index, by sampling the matrix from the tridiagonal matrix ensemble. However, as discussed, in this representation of the Gaussian $\beta$-ensemble, the ensemble with tridiagonal form is no longer invariant under symmetry transformations (unlike the classical Gaussian ensembles). In this section, we obtain the FS from a one-parameter family of $2\times 2$ random matrix ensembles introduced in [71], which is rotationally invariant but has an effective real value of the Dyson index ($\beta_{e}$). As we have shown in the following, when a matrix drawn from this ensemble is added to a diagonal integrable Hamiltonian, the resulting FS ones again diverges as one moves to the integrable phase, suggesting that this feature is generic for random matrix ensembles having an integrability-to-chaos transition, irrespective of whether the underlying ensemble is rotational invariant or not.
@@ VII.1 A $2\times 2$ matrix ensemble with an effective Dyson index
Consider the four random variables $u,v,x,y$ taken from a one-parameter ($\eta$) dependent joint probability distribution $\mathbb{P}_{\eta}(u,v,x,y)$ and an ensemble defined by the following Hermitian and unitary invariant $2\times 2$ matrices,
@@ VII.2 Fidelity susceptibility for $2\times 2$ invariant $\beta$-ensemble
We now use the $2\times 2$ matrix ensemble with an effective Dyson index to discuss the geometry of eigenstates when a matrix belonging to this ensemble is added to an integrable Hamiltonian. Specifically, the total Hamiltonian once again takes the form (1), where the integrable term $H_{0}$ is taken to be of the form
@@
Figure 9: Plot of $\tilde{g}_{rr}(r)$ for the effective $\beta$ ensemble with $\eta=1$, i.e., $\beta_{e}=0$ (black dots). The red curve shows a simple inverse square fitting of the numerical data: $\tilde{g}_{rr}(r)\approx 0.1/r^{2}$. The magenta dot-dashed curve is the plot for the $rr$ component of the line element for the GUE case (see eq. (55)). The plot in the inset shows more clearly the fitting for large values of $r$.
@@
Comparing plots with $\eta=1$ and $\eta=1/2$, another possible conclusion one can draw is that, even though there is a $(E_{m}-E_{n})^{2}$ term in the denominator of the definition of $\tilde{g}_{rr}(r)$, its decay with respect to $r$ may not be due to the correlation and repulsion between the eigenvalues but rather due to the fact that this energy difference term increases with increasing $r$. Thus, as $r\rightarrow\infty$, $\tilde{g}_{rr}(r)$ goes to zero even for $\eta=1$ the case where there is no correlation between the eigenvalues of the Hamiltonian of this invariant $2\times 2$ $\beta$-ensemble.
VIII Conclusions and Discussions
@@
In this paper we have considered the geometry of the quantum states for random matrix Hamiltonians that belong to some specific ensembles. The random matrix universality in the energy eigenvalues is commonly taken as a signature of chaotic behaviour of many-body quantum systems in the sense that fine-grained properties of the energy spectrum of these complex quantum many-body systems are well approximated by the random matrix theory [23, 24]. Nevertheless, it is also important to understand the properties of the eigenstates of these many-body Hamiltonians, possibly, as a first approximation, by modelling them with random matrices. Instead of using the conventional tools for this purpose, such as the inverse partition ratio of these eigenstates in a complete basis, here we have used the natural geometry of the space of quantum states induced by the Hermitian inner product, which measures the infinitesimal distance between two states. Here we focus specifically on random matrix Hamiltonians, which are drawn from ensembles that are more general than those of the three classical Gaussian ensembles and show a chaos-to-integrability transition. In particular, we have focused our attention on two generalisations of the classical Gaussian ensembles, namely the RP random matrix ensemble and the tridiagonal Gaussian $\beta$-ensembles, both of which show chaos-to-integrability transition in the level spacing distribution of the energy eigenvalues and break the rotational invariance of the classical Gaussian ensembles. [^17] We have also considered a specific form of the $\beta$-ensemble which keeps the invariance intact. It is possible to obtain the (suitably averaged) QMT on the parameter space (in some cases analytically) and explore how it encodes different characteristic properties of the eigenstates of the random matrix Hamiltonian under consideration.

### v2

@@ VI.1 Fidelity susceptibility for tridiagonal Gaussian $\beta$-ensembles
One difference from the situation considered in the two previous sections is that here we are considering the simplest situation where the total Hamiltonian $H$ has only one parameter $r$, rather than both $r$ and $\phi$. Hence, in this case, the line element will consist of a single element, which is the FS associated with the parameter $r$. The QMT components in such a situation are evaluated numerically in the next subsection.
For our purposes, it is sufficient to consider a system with a Hamiltonian of the form $H=H_{0}+r~\mathcal{H}_{\beta}$~ where $H_{0}$ is a diagonal ‘integrable’ Hamiltonian, i.e., we assume that its eigenvalues are drawn from independent Gaussian distributions with zero mean and variance unity, and $\mathcal{H}_{\beta}$~ is a ‘tridiagonal’ matrix. Therefore, we take $\mathcal{H}_{\beta}$ to be the following $2\times 2$ matrix
$\mathcal{H}_{\beta}=\left(\begin{array}[]{ccc}a_{0}&b_{1}\\ b_{1}&a_{1}\end{array}\right)~.$ (70)
@@
The Hamiltonian $H=H_{0}+r~\mathcal{H}_{\beta}$~ can be thought of as the generalisation of the usual RP model (which is usually defined with $\mathcal{H}$ taken from the classical Gaussian $\beta$-ensembles) for general values of $\beta$. The expression for $g_{rr}$, averaged over the ensemble is given by,
@@
Here $\mathcal{H}_{mn}=\braket{m|\mathcal{H}_{\beta}|n}$, are the matrix elements of $\mathcal{H}_{\beta}$ in the energy eigenbasis of the total Hamiltonian.
@@ VI.2 QMT for the tridiagonal $\beta$-ensemble
So far in this section, we have considered the simplest situation where the total Hamiltonian contains one parameter and the matrices are two-dimensional. Here, we now consider the more general situation where the total Hamiltonian is assumed to be of the form (57), where the $N\times N$ matrices $\mathcal{H}_{i}$ are now sampled from tridiagonal $\beta$-ensembles with generic Dyson indices. In Fig. 8 we have plotted the numerical values of the metric components for different fixed values of $r$ (and with a fixed $\phi$), when the matrix $H_{0}$ is drawn from a $\beta$-ensemble with a Dyson index $\gamma<1$, and $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are drawn from $\beta$-ensemble with the same Dyson index $\beta>1$. Note that, compared with the previous cases of GUE Hamiltonians, the mixed component $g_{r\phi}$ is non-zero, and the numerical values of this component are shown in the inset of Fig. 8.
@@
Figure 8: Log-Log (with base 10) plot of the components of the metric tensor for the Gaussian $\beta$-ensemble with large NN. The Hamiltonian is of the same form as (57), with the individual matrices being drawn from a $\beta$-ensemble with different Dyson indices ($H_{0}$ is from an ensemble with Dyson index $\gamma<1$, while $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are drawn from $\beta$-ensemble with the same Dyson index $\beta>1$). The red and the orange dots show the numerical values of the $g_{rr}$ and $g_{\phi\phi}$ components, respectively, for fixed values of $r$ and with $\beta=2$ and $\gamma=0.5$, while the brown and black dots represent these components when $\beta=3$ and $\gamma=0.2$. The plot in the inset shows the values of the $g_{r\phi}$ component for these two cases (magenta: $\beta=2$ and $\gamma=0.5$, and green: $\beta=3$ and $\gamma=0.2$). For all the plots, we have fixed $N=1000$, and the results show expressions averaged over 1000 different realisations of the ensemble.
@@ VII The fidelity-susceptibility for an invariant $\beta$-ensemble with an effective Dyson index
In the previous section we have obtained the FS associated with a $2\times 2$ random matrix ensemble with general Dyson index, by sampling the matrix from the tridiagonal $\beta$-ensemble. However, as discussed, in this representation of the Gaussian $\beta$-ensemble, the ensemble with tridiagonal form is no longer invariant under symmetry transformations (unlike the classical Gaussian ensembles). To see what effect this invariance has on the geometry of corresponding energy eigenstates, in this section, we compute the FS from a one-parameter family of $2\times 2$ random matrix ensembles introduced in [81], which is rotationally invariant but has an effective real value of the Dyson index ($\beta_{e}$). As we show in the following, when a matrix drawn from this ensemble is added to a diagonal integrable Hamiltonian, the resulting FS ones again diverges as one moves to the integrable phase, suggesting that this feature is generic for random matrix ensembles having an integrability-to-chaos transition, irrespective of whether the underlying ensemble is rotational invariant or not.
@@ VII.1 A $2\times 2$ matrix ensemble with an effective Dyson index
Consider four random variables $u,v,x,y$ taken from a one-parameter ($\eta$) dependent JPD $\mathbb{P}_{\eta}(u,v,x,y)$ and an ensemble defined by the following Hermitian and unitary invariant $2\times 2$ matrices,
@@ VII.2 Fidelity susceptibility for $2\times 2$ invariant $\beta$-ensemble
We now use this $2\times 2$ matrix ensemble with an effective Dyson index to obtain the geometry of eigenstates when a matrix belonging to this ensemble is added to an integrable Hamiltonian. Specifically, the total Hamiltonian once again takes the form (1), where the integrable term $H_{0}$ is taken to be of the form
@@
Figure 9: Plot of $\tilde{g}_{rr}(r)$ for the effective $\beta$-ensemble with $\eta=1$, i.e., $\beta_{e}=0$ (black dots). The red curve shows a simple inverse square fitting of the numerical data: $\tilde{g}_{rr}(r)\approx 0.1/r^{2}$. The magenta dot-dashed curve is the plot for the $rr$ component of the line element for the GUE case (see eq. (58)). The plot in the inset shows more clearly the fitting for large values of $r$.
@@
Comparing plots with $\eta=1$ and $\eta=1/2$, another possible conclusion one can draw is that, even though there is a $(E_{m}-E_{n})^{2}$ term in the denominator of the definition of $\tilde{g}_{rr}(r)$, its decay with respect to $r$ may not be due to the correlation and repulsion between the eigenvalues but rather due to the fact that this energy difference term increases with increasing $r$. Thus, as $r\rightarrow\infty$, $\tilde{g}_{rr}(r)$ goes to zero even for $\eta=1$ - the case where there is no correlation between the eigenvalues of the Hamiltonian of this invariant $2\times 2$ $\beta$-ensemble.
VIII Energy level dynamics: ensemble-averaged level curvature
VIII.1 Second derivative of the energy levels
Another interesting quantity, which is very similar to the FS considered so far, however, is a quantifier of the nature of the dependence of the energy levels of a complex many-body quantum system on the parameters (rather than the eigenstates), is the energy level curvature. For a Hamiltonian of the form in eq. (57), the curvature of the $n$th energy level $\mathcal{K}_{n}(r)$ is defined as the second derivative of $E_{n}(r)$ with respect to the parameter $r$, i.e., [^19] Compared to the standard definition of the level curvature that exists in the literature [51, 52, 49, 50, 82], we have absorbed a factor of 2 in the definition of $K_{n}$ for convenience. Also note that the expression in (86) is valid only for a Hamiltonian which can be written as $H(r)=H_{0}+r\mathcal{H}$. The general expression for the second-order derivative of $E_{n}(r)$ of a generic Hamiltonian $H(r)$ contains an additional term which is proportional to $\braket{n(r)|d^{2}H(r)/dr^{2}|n(r)}$ [49, 50]. Furthermore, the sum of energy level curvatures over all the eigenvalues $E_{n}(r)$ is trivially zero.
$\mathcal{K}_{n}=\frac{1}{2}\frac{d^{2}E_{n}(r)}{dr^{2}}=\sum_{j(\neq n)}\frac{\mathcal{H}_{nj}\mathcal{H}_{jn}}{(E_{n}(r)-E_{j}(r))}~.$ (86)
This can be rewritten as
$\mathcal{K}_{n}=-\int_{-\infty}^{\infty}~\frac{\text{d}\omega}{\omega}~\mathcal{S}_{n}(\omega)~,$ (87)
where the function $\mathcal{S}_{n}(\omega)$ defined in eq. (10). Comparing with the expression for the moments of $\mathcal{S}_{n}(\omega)$ in eq. (12), we see that moment with k=1k=-1 is just the $n$-th energy level curvature, i.e., $\mathcal{K}_{n}=-M_{-1}$.
Furthermore, defining the quantity $\chi_{\mathcal{H}}(z)$ through the Cauchy transformation of the function $\mathcal{S}_{n}(\omega)$,
$\chi_{\mathcal{H}}(z)=\int_{-\infty}^{\infty}~\frac{\text{d}\omega}{z-\omega}~\mathcal{S}_{n}(\omega)~,$ (88)
we see that the $n$-th level curvature can be written as $\mathcal{K}_{n}=\chi_{\mathcal{H}}(0)$. Note that the relation in (88) is analogous to the relation between the dynamical susceptibility (which is an appropriate extension of the advanced or retarded response function in the complex plane) and the spectral function of a Hermitian operator in the linear response theory [83, 84].
VIII.2 Energy level curvature for tridiagonal $\beta$-ensemble
Next, we compute the ensemble-averaged energy level curvature for an eigenstate of the Hamiltonian $H(r)=H_{0}+r\mathcal{H}$, with $H_{0}$ and $\mathcal{H}$ being the same matrices considered in section VI.1 (eqs. (71) and (70), respectively). Computation of the ensemble-averaged $\mathcal{K}_{n}(r)$ for a specific value $n$ proceeds along a similar line as the FS. It is possible to obtain an exact expression for $\bar{\mathcal{K}}_{2}(r,\beta)$ (or $\bar{\mathcal{K}}_{1}(r,\beta)$) for a generic value of the Dyson index $\beta>0$ for the $2\times 2$ matrix ensemble by performing the relevant integrations. However, the final expression for it is quite lengthy, so we omit it here. For specific values of $\beta$, the expressions do get simplified; e.g., with $\beta=2$ we get
$\bar{\mathcal{K}}_{2}(r,\beta=2)=\frac{1}{4\sqrt{\pi}}\bigg{(}\text{arccosh}\Big{(}\frac{4}{r^{2}}+1\Big{)}-\frac{2\sqrt{2}}{\sqrt{2+r^{2}}}\bigg{)}~,$ (89)
while for $\beta=4$ the expression simplifies to
$\bar{\mathcal{K}}_{2}(r,\beta=4)=\frac{1}{32\sqrt{\pi}}\bigg{(}(8+3r^{2})~\text{arccosh}\Big{(}\frac{8}{r^{2}}+1\Big{)}-12\sqrt{4+r^{2}}\bigg{)}~.$ (90)
Series expansions of these quantities for small values of $r$ show that they behave as $\log r$ for small values of $r$. While, for large values of $r$, these go to zero. In Fig. 11, we have shown the plots of $\bar{\mathcal{K}}_{2}(r,\beta)$ for $\beta=2,4$ along with the plot of the FS with $\beta=2$ derived in section VI.1 (see the expression of $\tilde{g}_{rr}(r)$ in eq. (58)).
Figure 11: Plot of the second energy level curvature $\bar{\mathcal{K}}_{2}(r,\beta)$ for the total Hamiltonian $H(r)=H_{0}+r\mathcal{H}$, where $H_{0}$ is a diagonal integrable matrix, and $\mathcal{H}$ is sampled from tridiagonal $\beta$-ensemble with generic $\beta$: $\beta=1$ (dashed brown) $\beta=2$ (red) and $\beta=4$ (dashed magenta). The eigen level curvature behaves as $\log r$ for small values $r$ and vanishes at large $r$ for generic values of $\beta>0$. For comparison, we have also plotted the ensemble-averaged FS obtained in the same scenario with $\beta=2$ (the green curve), which behaves as $1/r$ for small $r$.
VIII.3 Third derivative of the energy levels
Let us now consider the third derivative of the energy levels, which can be written, for a Hamiltonian of the form $H(r)=H_{0}+r\mathcal{H}$ as
$\begin{split}\mathcal{L}_{n}=\frac{1}{6}\frac{d^{3}E_{n}(r)}{dr^{3}}\hskip 85.35826pt\\ =\frac{1}{6}\Big{(}\braket{n^{\prime\prime}(r)|\mathcal{H}|n(r)}+\braket{n(r)|\mathcal{H}|n^{\prime\prime}(r)}+2\braket{n^{\prime}(r)|\mathcal{H}|n^{\prime}(r)}\Big{)}~.\end{split}$ (91)
Using the formula for the $\braket{n^{\prime\prime}(r)|k(r)}$ for $k\neq n$ obtained from the eigenvalue equation for $H(r)$, we get the following expression for the third derivative $\mathcal{L}_{n}$,
$\mathcal{L}_{n}(r)=\sum_{k(\neq n),m(\neq n)}\frac{\mathcal{H}_{kn}\mathcal{H}_{nm}\mathcal{H}_{mk}}{E_{nk}E_{nm}}-\mathcal{H}_{nn}\sum_{k(\neq n)}\frac{|\mathcal{H}_{nk}|^{2}}{E_{nk}^{2}}~,$ (92)
where $E_{nk}=E_{n}-E_{k}$. Note that the second term in the above expression is proportional to the FS of the $n$-th eigenstate, eq. (3).
We can obtain the connection with the correlation functions, as in the case of the FS and the second derivative of $E_{n}$, as follows. First note that since the second term above is proportional to the FS, it can be written in terms of the function $\mathcal{S}_{n}(\omega)$ (which is the Fourier transform of the connected two-point correlation function). Using a similar strategy to the manipulation of eq. (8), we can write the first term in (92) as
$\int\frac{\text{d}\omega}{\omega}\frac{\text{d}\omega^{\prime}}{\omega^{\prime}}\mathcal{Q}_{n}(\omega,\omega^{\prime})$~ (93)
where we have defined the function $\mathcal{Q}_{n}(\omega,\omega^{\prime})$, which is a successive Fourier transformation of a three-point correlation function:
$\mathcal{Q}_{n}(\omega,\omega^{\prime})=\int\frac{\text{d}t}{2\pi}\frac{\text{d}t^{\prime}}{2\pi}e^{-i(t\omega+t^{\prime}\omega^{\prime})}\Big{(}\braket{n|\mathcal{H}(t^{\prime})\mathcal{H}(0)\mathcal{H}(-t)|n}-2\braket{n|\mathcal{H}(t)\mathcal{H}(0)|n}\mathcal{H}_{nn}+\mathcal{H}_{nn}^{3}\Big{)}~.$ (94)
To ensure the convergence of these integrals, sometimes one has to necessarily introduce factors of $\exp({-\epsilon|t|})$. Hence, the final expression for $\mathcal{L}_{n}(r)$ in terms of the correlation functions is given by
$\mathcal{L}_{n}(r)=\int_{-\infty}^{\infty}\frac{\text{d}\omega}{\omega}\frac{\text{d}\omega^{\prime}}{\omega^{\prime}}\mathcal{Q}_{n}(\omega,\omega^{\prime})-\int_{-\infty}^{\infty}~\frac{\text{d}\omega}{\omega^{2}}~\mathcal{P}_{n}(\omega)~,$ (95)
where $\mathcal{P}_{n}(\omega)=\mathcal{H}_{nn}\mathcal{S}_{n}(\omega)$.
IX Conclusions and Discussions
@@
In this paper we have considered the geometry of the quantum states for random matrix Hamiltonians that belong to some specific ensembles. The random matrix universality in the energy eigenvalues is commonly taken as a signature of chaotic behaviour of many-body quantum systems in the sense that fine-grained properties of the energy spectrum of these complex quantum many-body systems are well approximated by the random matrix theory [24, 25]. Nevertheless, it is also important to understand the properties of the eigenstates of these many-body Hamiltonians, possibly, as a first approximation, by modelling them with random matrices. Instead of using the conventional tools for this purpose, such as the inverse partition ratio of these eigenstates in a complete basis, here we have used the natural geometry of the space of quantum states induced by the Hermitian inner product, which measures the infinitesimal distance between two states. Here, we focus specifically on random matrix Hamiltonians, which are drawn from ensembles that are more general than those of the three classical Gaussian ensembles and show a chaos-to-integrability transition. In particular, we have focused our attention on two generalisations of the classical Gaussian ensembles, namely the RP random matrix ensemble and the tridiagonal Gaussian $\beta$-ensembles, both of which show chaos-to-integrability transition in the level spacing distribution of the energy eigenvalues and break the rotational invariance of the classical Gaussian ensembles. [^20] We have also considered a specific form of the $\beta$-ensemble which keeps the invariance intact. It is possible to obtain the (suitably averaged) QMT on the parameter space (in some cases analytically) and explore how it encodes different characteristic properties of the eigenstates of the random matrix Hamiltonian under consideration.